OpenStudy (studydino1201):

A box contains 3 bags of corn chips, 3 bags of potato chips, 3 bags of peanuts, and 3 bags of pretzels. One bag is chosen at random. What is the probability that it is a bag of peanuts? A. 1/3 B. 1/4 C. 1/12 D. 1/2

OpenStudy (studydino1201):

@mathstudent55

OpenStudy (anthonyym):

How many bags of peanuts are there and how many total bags are there?

OpenStudy (mathstudent55):

\(\small probability(desired ~outcome) = \dfrac{number ~of ~ways ~desired ~outcome ~can ~happen}{total ~number ~of ~outcomes} \)

OpenStudy (mathstudent55):

The total number of outcomes is the sum of the numbers of all types of bags. The number of ways the desired outcome can happen is the number of bags of the type you want.

OpenStudy (studydino1201):

so C

OpenStudy (studydino1201):

or D

OpenStudy (mathstudent55):

\(\small p(peanuts) = \dfrac{3~ bags ~peanuts}{3 ~bags ~corn ~chips + 3 ~bags ~potato ~chips + 3 ~bags ~peanuts +3 ~bags ~pretzels} \)

OpenStudy (mathstudent55):

\(\small p(peanuts) = \dfrac{3~bags}{3 ~bags + 3 ~bags + 3 ~bags +3 ~bags } \) \(\small p(peanuts) = \dfrac{3~bags}{12 ~bags}\)

OpenStudy (mathstudent55):

What does the fraction 3/12 reduce to?

OpenStudy (studydino1201):

1/3

OpenStudy (mathstudent55):

No. To reduce a fraction, you must divide the numerator and denominator by the same number. \(\dfrac{3}{12} = \dfrac{3 \times 1}{3 \times 4} = \dfrac{3}{3} \times \dfrac{1}{4} = 1 \times \dfrac{1}{4} = \dfrac{1}{4}\)